OpenStudy (anonymous):

Can someone help me with this problem: Kyle is flying a helicopter and is rising at 5.0m/s when he releases the bag. After 2.0s a) what is the bag's velocity? b) how far has the bag fallen? c) how far below from the helicopter is the bag?

OpenStudy (anonymous):

a)\[v = u +at \] so , v = -5 +g*2 v = 14.6 b)\[s = ut + 1/2 a t^{2}\] s = -5 *2 + 1/2 *g * 2^2 s = 9.6 c) s (from heli) = 9.6 + 10 = 19.6